# Straightening the Ruler: Field-Level Inference of the BAO Scale with LEFTfield

### v1

@@ 6.1 BAO Reconstruction Procedure
We now describe how we implement a BAO reconstruction pipeline on our mock data. The initial step involves generating mock datasets using the LEFTfield code. The parameters chosen for the creation of these datasets are aligned with those employed in the field-level inference (as described below) to maintain consistency across the two analyses. This includes the use of the same bias parameters, noise levels, and the value for $\beta$. A flowchart summarizing both analyses is shown in Fig. 6.
We adopt the standard reconstruction procedure, which is based on the algorithm introduced by [5]. However, before this can be applied, we need to convert the density distribution on a grid (of size $N^{\Lambda_{0}}=192^{3}$, where $\Lambda_{0}=0.3\,h\,{\rm Mpc}^{-1}$ is the cutoff value used to generate the mock) generated by the LEFTfield code to a catalog of discrete “tracer” positions. This is done by first replacing the Gaussian likelihood used in the EFT likelihood with a Poisson process, which generates a discrete tracer count $N_{i}$ in each voxel $i$. For the average tracer number density $\bar{n}_{g}$, we adopt the value of $1/P_{\epsilon}$ used for the mocks employed in the field-level analysis. We then assign random positions to each of the $N_{i}$ tracers within the voxel $i$.
Figure 6: Diagram of two methods of BAO inference; field level inference (top) and power spectrum-based inference with and without reconstruction (bottom). The reconstruction step includes the generation of a discrete tracer catalog from $\delta_{g}$, as described in the text
Once the mock catalog is obtained, we proceed with the reconstruction. To ensure a fair comparison with the field level approach and guarantee that both methods have access to the same $k$-modes, we carefully select the smoothing scale and grid sizes. The choice of the smoothing scale is particularly important, given that the field-level approach employs a sharp-$k$ filter, while standard reconstruction involves a Gaussian filter. We choose the Gaussian smoothing scale $R=\frac{1}{\Lambda}$. We consider this to be a conservative choice for the comparison, as it allows for a significant contribution from modes with $k>\Lambda$ in the standard reconstruction approach, while these modes are excluded from the field-level analysis.
The reconstruction proceeds as follows.
@@
    Smooth the density field $\delta_{g}$ using a Gaussian filter $S(k,R)$ with a smoothing scale $R$: $\delta_{g}(\mathbf{k})\rightarrow S(k,R)\delta_{g}(\mathbf{k})$.
@@
    Using the smoothed density field, we find the estimated displacement $\boldsymbol{\psi}$, where
    $\boldsymbol{\psi}(\mathbf{k})\equiv-i\frac{\mathbf{k}}{k^{2}}S(k,R)\frac{\delta_{g}(\mathbf{k})}{b_{\delta}}.$ (6.1)
@@
    We interpolate $\boldsymbol{\psi}$ to find its value at the position of each tracer and use it to move the tracers.
@@
    Once all the tracers have been shifted, we use the nearest-grid-point (NGP) assignment scheme and a grid of size $N^{\Lambda_{0}}$ to obtain the “displaced” density field $\delta_{d}$. Notice that this operation removes a large fraction of the large-scale perturbations in in $\delta_{g}$.
@@
    We generate a spatially uniform grid of particles and shift them by $\boldsymbol{\psi}$ to create the “shifted” field $\delta_{s}$. The assignment scheme used to find $\delta_{s}$ is again NGP, and the grid size is $N^{\Lambda_{0}}$. Notice that this operation removes a large fraction of the large-scale perturbations in in $\delta_{g}$.
@@
  7. 7.
    Next, we resize the reconstructed field to a grid of size $N_{\text{g}}^{\Lambda}$ to make sure we keep only the modes below the cutoff.
@@
As an illustration, the left panel of Fig. 7 displays the oscillatory part of the power spectrum averaged over 500 Mock A-like realizations, for both pre- and post-reconstruction cases. Post-reconstruction, the wiggles become significantly more pronounced.
@@
Figure 7: The left panel shows the oscillatory part of the power spectrum averaged over 500 Mock A-like realizations. We show this for both the pre- and post-reconstructed power spectrum (for $\Lambda=0.2\,h\,{\rm Mpc}^{-1}$). We see that following the reconstruction, the wiggles on smaller scales become more pronounced. The right panel shows the inferred error bar $\sigma_{\text{PS}}(\beta)$ on the BAO scale using pre- and post-reconstruction power spectrum. The gain in BAO scale precision due to reconstruction increases as $\Lambda$ increases.
Finally, to obtain the BAO scale parameter $\beta$ and its uncertainty, we use a maximum-likelihood approach, where all parameters apart from $\beta$ are fixed to their ground truths. For a given value of $\Lambda$, we calculate the power spectrum likelihood $\mathcal{L}_{\text{PS}}$ for a set of $\{\beta_{i}\}$ values, where
@@
$-2\log\mathcal{L}_{\text{PS}}(\beta)=\frac{\left(P_{\text{data}}(k)-P_{\text{theory}}(k,\beta)\right)^{2}}{\text{Cov}[P_{\text{data}}(k)]}+{\rm const}.$ (6.2)
Here, $P_{\text{data}}(k)$ represents the power spectrum of the reconstructed mock data from which we aim to infer the BAO scale, and $P_{\text{theory}}(k,\beta)$ is the theoretical prediction for the reconstructed power spectrum for different $\beta$ values. We find the latter by averaging over 500 samples, each sharing the same cosmological, bias, and noise parameters at fixed $\beta$, but featuring independent realizations of the initial conditions. We use linear binning to find the power spectrum. The forward model, specified by $\Lambda_{0}$, as well as all bias terms and $\sigma_{\epsilon}$ are fixed at the values used for generating the mock. It is worth noting that $P_{\text{theory}}(k,\beta)$ already incorporates the stochastic contributions $P_{\epsilon}$. We use the linear covariance
$\text{Cov}[P_{\text{data}}(k)]=2\frac{\left[P_{\rm theory}(k,\beta_{0})\right]^{2}}{m_{k}},$ (6.3)
where $m_{k}$ is the number of modes inside a bin. Since the covariance is independent of $\beta$, the normalization constant in Eq. (6.2) is irrelevant.
We then fit a parabola to the set of $\{\beta_{i},-2\ln\mathcal{L}_{\text{PS}}(\beta_{i})\}$. The maximum-likelihood value of $\beta$, denoted as $\hat{\beta}$, is located at the minimum of this parabola. The error bar, $\sigma_{\rm PS}(\beta)$, can be estimated by finding the point at which $-2\log\mathcal{L}_{\text{PS}}$ increases by 1. Note that we apply this inference procedure using $\mathcal{L}_{\text{PS}}$ both for the pre- and post-reconstruction power-spectra, as depicted in Fig. 6. The improvement in BAO scale determination after reconstruction is demonstrated in the right panel Fig. 7. Notice that the relative improvement grows with with increasing $\Lambda$, which is expected since the sharpening of the BAO by reconstruction is more pronounced at higher $k$.
Compared to the field-level approach, which jointly infers all bias and stochastic parameters, this analysis strongly favors the standard power-spectrum-based method. By conducting the reconstruction at the same cutoff used to generate the mock, we give access to the complete information about the mock, and avoid the model mismatches that the field-level analysis had to overcome. Additionally, we fix the values of the bias parameters, whereas in practice, jointly inferring the values of the bias parameters would introduce further uncertainties into the inference process and enlarge the error bars. These issues are clearly worth revisiting in the future.
@@ 6.2 Inference results and comparison with the field level
In this section, we compare the results of the power spectrum analysis (before and after reconstruction) with the field-level inference presented in the previous section.
We begin with Mock A, presenting in Fig. 8 the ratio of error bars on $\hat{\beta}$ derived from the two power-spectrum-based analyses, $\sigma_{\text{PS}}$, to those obtained at the field level, $\sigma_{\text{F}}$. Compared to the field level, the pre-reconstruction power spectrum yields an error bar between 1.15 and 2 times larger. As expected, this difference between $\sigma_{\text{PS}}$ and $\sigma_{\text{F}}$ decreases once we perform reconstruction, where we find $\sigma_{\text{PS}}$ to be at most 1.5 times larger than $\sigma_{\text{F}}$. Note that the improvemend of field level over power spectrum generally increases with the cutoff. There are several possible explanations for this, which we turn to in the discussion Sec. 7.
Figure 8: Inferred error bar $\sigma_{\text{PS}}(\beta)$ on the BAO scale using pre- and post-reconstruction power spectrum, relative to that in the field-level inference $\sigma_{\text{F}}$, in the case of Mock A. This mock was generated (and sampled, in case of field-level inference) using Lagrangian bias. Pre-reconstruction results are depicted using squares, whereas post-reconstruction results are represented by stars. It is evident that, even comparing to power spectrum after BAO reconstruction, the field-level BAO scale inference is more precise, by up to a factor of 1.4, corresponding to a $\sim 2$ times larger survey volume.
Figure 9: Inferred error bar $\sigma_{\text{PS}}(\beta)$ on the BAO scale from the post-reconstruction power spectrum, relative to that in the field-level inference $\sigma_{\text{F}}$, in the case of Mock B. Blue points indicate the scenario where the Lagrangian bias is used in the field level inference, while orange points correspond to the Eulerian case (the reconstruction results are the same in both cases). The field-level error bar is consistently smaller, regardless of which bias basis is used in the field-level inference.
We perform a similar analysis in the case of the Mock B, the results of which are summarized in Fig. 9. In the plot, the orange points represent the comparison of the post-reconstruction power spectrum variance $\sigma_{\text{PS}}$ to the case when we use Eulerian bias expansion in the field level. Blue points correspond to Lagrangian bias expansion being used for the field level. We find that results here are consistent with the ones in the case of Mock A: $\sigma_{\text{PS}}$ (after reconstruction) is between 1.1 and 1.6 times larger than $\sigma_{\text{F}}$, with the improvement generally increasing toward higher $\Lambda$.
@@ 7 Discussion and Conclusions
In this paper, we have presented the outcome of field-level BAO scale inferences, which jointly infer the BAO scale together with the initial conditions, bias and stochastic parameters via the LEFTfield code, applied to mock data in the rest frame (without redshift-space distortions). The mock datasets were created using bias parameters taken from a fixed initial condition (FixedIC) analysis on halo catalogues, and at a substantially higher cutoff (or resolution) than those used in the inference. By introducing model misspecification in this way, we attempt to make our mocks as realistic as possible. The primary distinction between the two sets of mock data is in the bias expansion model employed for their generation: Mock A was produced using the second-order Lagrangian bias, whereas Mock B used the second-order Eulerian bias.
@@
For Mock A, the analysis was done only using the second-order Lagrangian bias expansion in the inference. It showed that the systematic bias in $\beta$ stays below 1% and is effectively negligible for all $\Lambda$ values, except for the highest cutoff, where it increases to about 1.8%. We plan to investigate this increase in systematic bias in $\beta$ in the future by going to the third order in the bias expansion.
@@
For the Eulerian-bias Mock B, our analysis included both second-order Lagrangian and Eulerian biases. For the Eulerian bias, we find that the remaining systematic bias is consistent with zero and decreases as $\Lambda$ increases. In the case of Lagrangian bias expansion, the remaining systematic bias is slightly higher. Again, it is possible that the residual systematic shift in the Lagrangian analysis decreases once we go to a higher order in bias expansion. Furthermore, we find that the error bar in the case of the Lagrangian analysis is lower than in the case of Eulerian analysis. This is likely connected to the approaches used for constructing bias operators in these models. For the Eulerian bias model, the construction of bias operators involves an additional filtering of the matter field, beyond the initial condition filtering. This extra filtering step results in the omission of some mode-coupling contributions, which are preserved in the Lagrangian bias model. The latter only applies filtering to the initial conditions. Note that both models utilize the same likelihood filter.
@@
The standard approach to BAO inference is based on applying a reconstruction procedure to the tracer catalog, and then measuring the BAO scale in the post-reconstruction power spectrum. In Sec. 6, we detailed the application of this reconstruction algorithm to our mock data. Throughout the reconstruction process, we employed the same $\Lambda_{0}$, and fixed the bias parameters to the values that were used to generate the mock datasets, thereby giving a significant benefit to the reconstruction-based analysis pipeline. Figures 8 and 9 summarize the comparison between the error bars obtained from the field level analysis, $\sigma_{\text{F}}$, and those derived using the power spectrum, $\sigma_{\text{PS}}$. These figures illustrate that, depending on the mock and the value of $\Lambda$, $\sigma_{\text{F}}$ is smaller than $\sigma_{\text{PS}}$ by up to a factor of two (pre-reconstruction), or between 1.4 and 1.6 (post-reconstruction), the latter corresponding to an effective doubling of the survey volume. Additionally, we observe that the improvement over the power spectrum analysis increases as more modes are included with increasing $\Lambda$.
@@
These results are clearly very encouraging for the field-level approach. However, one might naturally ask two questions about these findings: first, where does the additional information come from? As argued in [15], the field-level approach extracts the BAO feature not only in the power spectrum, but also in higher $n$-point functions of the tracer, such as the bispectrum, as it does not rely on any compression of the data. In addition, our use of a 2LPT forward model means that the change of the local BAO scale in the presence of large-scale density perturbations (due to the different local expansion in the separate-universe picture) is consistently captured. On the other hand, the Zeldovich approximation used in Eq. (6.1) does not describe this effect correctly [4]. Both of these effects are expected to become more significant as smaller scales are included in the analysis, which could explain the increasing improvement that we find toward smaller scales (higher $\Lambda$).
@@
It is worth noting that the maximum wavenumbers included in our analysis are somewhat lower than those chosen in current BAO reconstruction analyses. Specifically, these choose a smoothing kernel in the displacement construction on similar scales as the one used here, but then measure the post-reconstruction power spectrum to higher $k\sim 0.3-0.5\,h\,{\rm Mpc}^{-1}$. This approach assumes that the BAO are still well described by the perturbative model on scales where perturbation theory no longer applies, an assumption which needs to be carefully validated. On the other hand, the field-level approach by construction requires that the model describe all aspects of the data up to the maximum scale included, which precludes us from pushing this approach beyond perturbative scales.
@@
The second natural question is whether these results apply to actual nonlinear tracers. Thus, our next step will be to transition from mock data to dark matter halos in full N-body simulations, and perform a similar comparison there. We are also advancing our redshift-space modeling [25], which will enable an anisotropic BAO scale inference in terms of line-of-sight and perpendicular components. Moreover, we plan to jointly infer the BAO scale while varying cosmological parameters governing the growth, such as $\sigma_{8}$ and the growth rate $f$.
@@
Finally, the field-level forward model also enables a consistent Bayesian inference of the BAO scale using the BAO reconstruction technique, by way of simulation-based inference (SBI) [46], allowing us to consistently marginalize over bias, noise and cosmological parameters. We also plan to explore this direction in the near future.

### v2

@@ 6.1 BAO Reconstruction Procedure
We now describe how we implement a BAO reconstruction pipeline on our mock data. We follow a standard reconstruction procedure based on the algorithm introduced by [5]. However, in order for this to be applied, we need the mock data generated by the LEFTfield code in the form of a catalog of discrete “tracer” positions. This catalog is constructed as follows.
We first generate a noise-free mock catalog, using the same forward model, bias parameter values, and $\beta$ values as used for the mock catalogs described in Sec. 4. This yields a density distribution on a grid of size $N^{\Lambda_{0}}=192^{3}$, where $\Lambda_{0}=0.3\,h\,{\rm Mpc}^{-1}$ is the cutoff value used to generate the mock. We then Poisson sample a discrete tracer count $N_{i}$ in each voxel $i$ of the grid, where the voxel-dependent expectation value is $\langle N_{i}\rangle=\bar{N}(1+\delta_{\rm mock,noise-free})$. $\bar{N}=\bar{n}_{g}V_{\rm voxel}$ is determined via the mean tracer number density $\bar{n}_{g}$, which we set to $\bar{n}_{g}=1/P_{\epsilon}$, using the same $P_{\epsilon}$ as in the generation of the mocks introduced in Sec. 4. Finally, we assign random positions to each of the $N_{i}$ tracers within their respective voxels.
Figure 6: Flowchart of the two methods for BAO inference employed in this work: field-level inference (top) and power spectrum-based inference with and without reconstruction (bottom). The reconstruction step includes the generation of a discrete tracer catalog from $\delta_{g}$, as described in the text.
Once the mock catalog is obtained, we proceed with the reconstruction. To ensure a fair comparison with the field-level approach and guarantee that both methods have access to the same $k$-modes, we carefully select the smoothing scale and grid sizes. The choice of the smoothing scale is particularly important, given that the field-level approach employs a sharp-$k$ filter, while standard reconstruction involves a Gaussian filter. We choose the Gaussian smoothing scale $R=\Lambda^{-1}$. We consider this to be a conservative choice for the comparison, as it allows for a significant contribution from modes with $k>\Lambda$ in the standard reconstruction approach, while these modes are excluded from the field-level analysis. The reconstruction for a given cutoff $\Lambda<\Lambda_{0}$ proceeds as follows (see Fig. 6 for a flowchart summarizing both analyses):
@@
    The tracers are assigned to a grid of size $N_{\rm assign}=2N^{\Lambda}$ using the NUFFT assignment scheme to obtain the initial tracer field $\delta_{g}$.
@@
    The density field $\delta_{g}$ is smoothed using a Gaussian filter $W_{\rm G}(kR)$ with a smoothing scale $R=1/\Lambda$: $\delta_{g}(\bm{k})\rightarrow W_{\rm G}(k/\Lambda)\delta_{g}(\bm{k})$.
@@
    Using the smoothed density field, we find the estimated displacement $\bm{\psi}$, defined as
    $\bm{\psi}(\bm{k})\equiv-i\frac{\bm{k}}{k^{2}}W_{\rm G}(k/\Lambda)\frac{\delta_{g}(\bm{k})}{b_{\delta}},$ (6.1)
    where we employ the ground-truth value for $b_{\delta}$ used to generate the mock.
@@
    We interpolate $\bm{\psi}$ to find its value at the position of each tracer and use it to move the tracers.
@@
    Once all the tracers have been shifted, we use the NUFFT assignment scheme and a grid of size $2N^{\Lambda}$ to obtain the “displaced” density field $\delta_{d}$. Notice that this operation removes a large fraction of the large-scale perturbations in $\delta_{g}$.
@@
    We generate a spatially uniform grid of particles and shift them by $\bm{\psi}$ to create the “shifted” field $\delta_{s}$. The assignment scheme used to find $\delta_{s}$ is again NUFFT, and the grid size is $2N^{\Lambda}$.
  7. 7.
@@
As an illustration, the left panel of Fig. 7 displays the oscillatory part of the power spectrum averaged over 1000 Mock B-like realizations, for both pre- and post-reconstruction cases. Post-reconstruction, the wiggles become significantly more pronounced.
@@
Figure 7: The left panel shows the oscillatory part of the power spectrum averaged over 1000 Mock B-like realizations. We show this for both the pre- and post-reconstructed power spectrum (for $\Lambda=0.2\,h\,{\rm Mpc}^{-1}$). As expected, the wiggles become more pronounced following the reconstruction. The right panel shows the inferred error bar $\sigma_{\text{PS}}(\beta)$ on the BAO scale using the power spectrum of the data pre-reconstruction, relative to that from the post-reconstruction power spectrum, $\sigma^{\text{rec}}_{\text{PS}}(\beta)$. Both inferences use the likelihood given in Eq. (6.5).
We determine the BAO scale parameter $\beta$ and its uncertainty in the both pre- and post-recontruction analyses by performing an MCMC analysis using the emcee sampler [46]. The analysis jointly infers the BAO scale with broad-band power spectrum fitting parameters, using the following template:
$P_{\rm model}(k,\beta)=(B_{1}+B_{2}k^{2})P_{m}(k,\beta)+A(k),$ (6.2)
where $B_{1}$ and $B_{2}$ are fitting parameters, and
$P_{m}(k,\beta)=P_{\rm fid}(k)f^{2}(\beta,k)$ (6.3)
represents the linear power spectrum with the rescaled BAO feature, as introduced in Eq. (2.5). The term $A(k)$ is modeled as a third-order polynomial:
$A(k)=a_{0}+a_{2}k^{2}+a_{3}k^{3}.$ (6.4)
To compute the power spectrum, we apply linear binning and perform the fit within the range $0.03\,\,h\,{\rm Mpc}^{-1}<k<\Lambda$, with $\Delta k=0.0028\,h\,{\rm Mpc}^{-1}$.
The inference uses the following likelihood:
@@
$-2\log\mathcal{L}_{\text{PS}}(\beta)=\sum_{k_{i}}\frac{\left(P_{\text{data}}(k_{i})-P_{\text{model}}(k_{i},\beta)\right)^{2}}{\text{Cov}[P_{\text{best-fit}}(k_{i})]}+{\rm const},$ (6.5)
where the sum is over the linear bins in wavenumber with central values $k_{i}$, and $P_{\text{data}}(k)$ represents the power spectrum of the mock data from which we aim to infer the BAO scale. Depending on the case considered, this can be either the pre– or post-reconstruction power spectrum. It is worth noting that $P_{\text{model}}(k,\beta)$ already incorporates the stochastic contributions $P_{\epsilon}$ via the parameter $a_{0}$. We use the Gaussian covariance
$\text{Cov}[P_{\text{best-fit}}(k_{i})]=2\frac{\left[P_{\text{best-fit}}(k_{i},\beta_{0})\right]^{2}}{m_{i}},$ (6.6)
where $m_{i}$ is the number of modes inside the bin $i$. Before the MCMC analysis , the best-fit power spectrum, $P_{\text{best-fit}}$, is obtained by applying a curve-fitting procedure to determine the best-fit parameter values for the model function given the data $P_{\rm{data}}$. We stress that these parameters are only used to determine the covariance in the likelihood Eq. (6.5), which is then kept fixed. Since the covariance is constant (independent of all parameters that are varied), the normalization constant in Eq. (6.5) is irrelevant. For the parameter $\beta$, we adopt a uniform prior, $\mathcal{P}(\beta)=\mathcal{U}(0.6,1.4)$. For the remaining parameters, we use broad Gaussian priors with means set to the values obtained in $P_{\text{best-fit}}$. We have also found that the width of these priors does not affect the $\beta$ posterior.
For each $\Lambda$ value, we run 32 chains with 6000 samples, which gives us around 2800 effective samples. The improvement in BAO scale determination after reconstruction is demonstrated in the right panel of Fig. 7. Notice that the relative improvement grows with with increasing $\Lambda$, which is expected since the higher-$k$ wiggles are more affected by damping, and since the smaller smoothing scale employed increases the quality of the reconstruction.
@@ 6.2 Inference results and comparison with the field level
We are now ready to compare the results of the power spectrum analysis (before and after reconstruction) with the field-level inference presented in the previous section.
We begin with Mock A, presenting in Fig. 8 the ratio of error bars on $\hat{\beta}$ derived from the two power-spectrum-based analyses, $\sigma_{\text{PS}}$, to those obtained at the field level, $\sigma_{\text{F}}$. Compared to the field level, the pre-reconstruction power spectrum yields an error bar between 1.25 and 2 times larger. As expected, this difference between $\sigma_{\text{PS}}$ and $\sigma_{\text{F}}$ decreases once we perform reconstruction. At the lowest cutoff considered, $\Lambda=0.15\,h\,{\rm Mpc}^{-1}$, field-level inference and post-reconstruction power spectrum yield identical posteriors. For the higher cutoff values, the post-reconstruction error is $1.25-1.35$ times larger than $\sigma_{\text{F}}$. That is, field-level inference improves the precision of the BAO scale by $\sim 30\%$ over standard reconstruction.
Figure 8: Inferred error bar $\sigma_{\text{PS}}(\beta)$ on the BAO scale using pre- and post-reconstruction power spectrum, relative to that in the field-level inference $\sigma_{\text{F}}$, in the case of Mock A. This mock was generated (and sampled, in case of field-level inference) using Lagrangian bias. Pre-reconstruction results are depicted using squares, whereas post-reconstruction results are represented by stars. It is evident that, even comparing to power spectrum after BAO reconstruction, the field-level BAO scale inference is more precise, by up to a factor of 1.35.
Figure 9: Inferred error bar $\sigma_{\text{PS}}(\beta)$ on the BAO scale from the post-reconstruction power spectrum, relative to that in the field-level inference $\sigma_{\text{F}}$, in the case of Mock B. Blue points indicate the scenario where Lagrangian bias is used in the field level inference, while orange points correspond to the Eulerian case (the reconstruction results are the same in both cases).
We perform a similar analysis in the case of the Mock B, the results of which are summarized in Fig. 9. In the plot, the orange points represent the comparison of the post-reconstruction power spectrum variance $\sigma_{\text{PS}}$ to the case when we use the Eulerian bias expansion in the field-level inference, while blue points correspond to inferences based on the Lagrangian bias expansion (the post-reconstruction error bars are the same in both cases). These results show that the precision of the field-level inference of the BAO scale does depend on the forward model used. In this case, Eulerian bias yields an improvement also at the lowest cutoff, while at higher cutoffs Lagrangian bias leads to a better performance (we will discuss this in the following section). The comparison of Fig. 8 and Fig. 9 also shows that, for a given forward model used in the inference, the information gain in field-level BAO inference also depends on the dataset. Analyses of more independent mock catalogs will be necessary to definitely assess and understand both of these trends.
@@ 7 Discussion and Conclusions
In this paper, we have presented the outcome of field-level BAO scale inferences, which jointly infer the BAO scale together with the initial conditions, bias and stochastic parameters via the LEFTfield code, applied to mock data in the rest frame (without redshift-space distortions). The mock datasets were created using bias parameters taken from a fixed-initial-condition analysis on halo catalogues, and at a substantially higher cutoff (or resolution) than those used in the inference. By introducing model misspecification in this way, we attempt to make our mocks as realistic as possible. The primary distinction between the two sets of mock data is in the bias model employed for their generation: Mock A was produced using the second-order Lagrangian bias, whereas Mock B used the second-order Eulerian bias.
@@
For Mock A, the analysis was done only using the second-order Lagrangian bias expansion in the inference. It showed that the systematic bias in $\beta$ stays below 1% and is effectively negligible for all $\Lambda$ values, except for the highest cutoff, where it increases to about 1.8%. We have found that this bias is likely removed when moving to a third-order bias expansion in the inference.
@@
For the Eulerian-bias Mock B, our analysis included both second-order Lagrangian and Eulerian biases. For the Eulerian bias, we find that the remaining systematic bias is consistent with zero and decreases with increasing $\Lambda$. In the case of Lagrangian bias expansion, the remaining systematic bias is slightly higher. Again, it is possible that the residual systematic shift in the Lagrangian analysis decreases once we use a higher-order bias expansion. Furthermore, we find that the error bar in the case of the Lagrangian analysis is lower than in the case of Eulerian analysis. This is likely due to the additional filtering employed in the construction of Eulerian bias operators. This extra filtering step results in the omission of some mode-coupling contributions, which are preserved in the Lagrangian bias model.
@@
The standard approach to BAO inference is based on applying a reconstruction procedure to the tracer catalog, and then measuring the BAO scale in the post-reconstruction power spectrum. In Sec. 6, we detailed the application of this reconstruction algorithm to our mock data. We fixed the linear bias coefficient used in the reconstruction process, as is commonly done, thereby giving a slight benefit to the reconstruction-based analysis pipeline over field-level inference, where all bias coefficients are jointly inferred with the BAO scale. Figures 8 and 9 summarize the comparison between the error bars obtained from the field level analysis, $\sigma_{\text{F}}$, and those derived using the power spectrum, $\sigma_{\text{PS}}$. These figures illustrate that, depending on the mock dataset and the value of $\Lambda$, $\sigma_{\text{F}}$ is smaller than $\sigma_{\text{PS}}$ by up to a factor of two (pre-reconstruction), or up to 1.5 (post-reconstruction). Regardless of this difference, the field-level approach offers a more consistent framework for BAO inference, as it allows for the joint sampling of all parameters. Additionally, the field-level method does not require extra fitting templates such as Eq. (6.2) needed in the power spectrum analysis.
Our findings are further supported by the results of Sec. 5.2.1 and 5.2.2, where we show that the error bars on $\beta$ obtained via field-level inference precisely match those from a power-spectrum-based analysis within the context of a linear forward and bias model. Furthermore, when performing inference on a mock generated with a 1LPT (Zel’dovich) forward model with linear bias, using the same model in the inference, the error bars from field-level inference show excellent agreement with those obtained from the post-reconstruction power spectrum, which should be optimal for this specific setup. This latter test provides a strong, nontrivial validation of both the field-level inference approach and the BAO reconstruction procedure used throughout this work.
@@
These results are clearly very encouraging for the field-level approach. However, one might naturally ask two questions about these findings: first, where does the additional information come from? And second, do these results apply to actual nonlinear tracers? Part of the answer to the first question is the fact that field-level inference not only removes the large-scale displacements, but also nonlinear bias contributions in the forward evolution. Further, our use of a 2LPT forward model means that the change of the local BAO scale in the presence of large-scale density perturbations (due to the different local expansion in the separate-universe picture) is consistently captured. On the other hand, the Zel’dovich approximation used in Eq. (6.1) does not describe this effect correctly [4]. Both of these effects are expected to become more significant as smaller scales are included in the analysis, which could explain the increasing improvement that we find toward smaller scales (higher $\Lambda$). We will investigate the source of information gain in more detail in upcoming work.
@@
Throughout this paper, we have attempted to make a comparison between field-level inference and BAO reconstruction on a closely matched range of scales. It is worth noting that the maximum wavenumbers included in our analysis are somewhat lower than those chosen in current BAO reconstruction analyses. Specifically, these choose a smoothing kernel in the displacement construction on similar scales as the one used here, but then measure the post-reconstruction power spectrum up to $k_{\rm max}\sim[0.3-0.5]\,h\,{\rm Mpc}^{-1}$. This analysis assumes that the BAO are still well described by the perturbative model on scales where perturbation theory no longer applies, an assumption which needs to be carefully validated. On the other hand, the field-level approach by construction requires that the model describe all aspects of the data up to the maximum scale included, which precludes us from pushing this approach beyond perturbative scales.
@@
To address the second question, our next step will be to transition from mock data to dark matter halos in full N-body simulations, and perform a similar comparison there. The redshift-space modeling recently presented in [26, 47] will enable an anisotropic BAO scale inference in terms of line-of-sight and perpendicular components. Moreover, we plan to jointly infer the BAO scale while varying cosmological parameters governing the growth, such as $\sigma_{8}$ and the growth rate $f$.
@@
Finally, the field-level forward model also enables coupling the standard BAO reconstruction approach to a consistent Bayesian inference of the BAO scale, by way of simulation-based inference (SBI) [48], allowing us to consistently marginalize over bias, noise and cosmological parameters. We also plan to explore this direction in the near future.